# {textblock*}

### v2

@@ III.2 Fit to CMB and LSS data
In Fig. 3 we show the dependence of the 95 percentile of $f_{\mathrm{iso}}^{2}=(P_{S_{\gamma\nu}}+P_{S_{\gamma\mathrm{DM}}})/P_{\zeta}|_{k_{\mathrm{piv}}}$ as a function of $\varphi$  in orange. For comparison the bounds found in Ref. [8] are indicated by green arrows. As the analysis of Ref. [8] allowed the isocurvature spectral index to vary, we compare our results to the most stringent one out of the three scales considered in Tab. 14 of Ref. [8], the one at $k_{\mathrm{low}}\,$. More specifically, we took the values $f_{\mathrm{iso}}^{2}=2.5\%$ and $7.4\%$ for the dark matter and compensated neutrino case, respectively, where for the latter one we accounted for the difference in normalization by multiplying with $(1.4)^{2}$. Our results agree with Ref. [8] to within 30%, where the discrepancies are most likely due to differences in analysis. We further agree with Ref. [8] that the bounds are significantly more stringent when considering large correlations between the adiabatic and isocurvature mode (see Appendix C).
We find that the bound on the isocurvature fraction is weaker for scenarios with vanishing matter isocurvature $\varphi\approx\pi/2\sim-\pi/2$ This may be contributed to the fact that the signatures in the case of $S_{\gamma\mathrm{DM}}=0$ are suppressed on all scales relative to the adiabatic ones as can be seen from Fig. 2 (see also Appendix B). Furthermore, we find that the fit shows a slight preference for non-vanishing isocurvature for these values of $\varphi$ as the reduction of $\chi^{2}$ compared to a fit of a $\Lambda$CDM cosmology with varying $N_{\mathrm{eff}}$ exceeds the number $\Delta n=2$ of additional fit parameters ($f_{\mathrm{iso}}$, $c_{\zeta S}$).
@@ Acknowledgements.
We would like to thank J. Lesgourgues for insights and discussions on MontePython. The authors gratefully acknowledge the computing time granted on the supercomputer MOGON 2 at Johannes Gutenberg University Mainz (hpc.uni-mainz.de). PS and CG acknowledge support by the Cluster of Excellence “Precision Physics, Fundamental Interactions, and Structure of Matter” (PRISMA+ EXC 2118/1) funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) within the German Excellence Strategy (Project No. 390831469).
@@ Appendix A Initial conditions in coordinates
$h$ $-R_{\nu}\omega$ $\frac{3}{8}R_{\nu}\omega^{2}$ $\frac{R_{\nu}k^{2}\omega}{90}(3R_{\mathrm{b}}+4)-\frac{R_{\nu}\omega^{3}}{8}$
@@
$-\frac{R_{\nu}(16R_{\nu}(3R_{\nu}+40)-175)k^{2}\omega}{360(2R_{\nu}+15)(4R_{\nu}+15)}$
$\delta_{\gamma}$ $-\frac{4R_{\nu}}{3}$ $\frac{2R_{\nu}\omega\tau}{3}$ $\frac{1}{36}R_{\nu}\left(8k^{2}-9\omega^{2}\right)$
@@
$-\frac{2R_{\nu}(R_{\mathrm{b}}(R_{\nu}-6))k^{2}\omega}{90(R_{\nu}-1)}$
$\theta_{\gamma\mathrm{b}}/k$ $-\frac{R_{\nu}k}{3}$ $-\frac{R_{\nu}R_{\mathrm{b}}k\omega}{4(R_{\nu}-1)+\frac{R_{\nu}k\omega}{12}}$
@@
$-\frac{3R_{\nu}R_{\mathrm{b}}^{2}k\omega^{2}}{16(R_{\nu}-1)^{2}}$
@@
$-\frac{R_{\nu}k\omega^{2}}{48}$
$\sigma_{\nu_{\mathrm{eff}}}$ $-\frac{2(R_{\nu}-1)k^{2}}{12R_{\nu}+45}$ $\frac{(61-4R_{\nu})R_{\nu}k^{2}\omega}{18(2R_{\nu}+15)(4R_{\nu}+15)}$
$F_{3{\nu_{\mathrm{eff}}}}$ $-\frac{4(R_{\nu}-1)k^{3}}{21(4R_{\nu}+15)}$
$\delta_{\mathrm{b}}$ $-R_{\nu}$ $\frac{R_{\nu}\omega}{2}$ $\frac{1}{48}R_{\nu}\left(8k^{2}-9\omega^{2}\right)$
@@
$-\frac{R_{\nu}R_{\mathrm{b}}(R_{\nu}-6)k^{2}\omega}{60(R_{\nu}-1)}$
$\delta_{\mathrm{DM}}$ $-R_{\nu}$ $\frac{R_{\nu}\omega}{2}$ $-\frac{3}{16}R_{\nu}\omega^{2}$ $\frac{R_{\nu}\omega^{3}}{16}-\frac{R_{\nu}k^{2}\omega(3R_{\mathrm{b}}+4)}{180}$
@@
$h$ $R_{\mathrm{DM}}\omega$ $-\frac{3\omega^{2}}{8}R_{\mathrm{DM}}$ $\frac{R_{\mathrm{DM}}}{360}(45\omega^{3}-16k^{2}\omega)$
$\eta$ $-\frac{1}{6}R_{\mathrm{DM}}\omega$ $\frac{\omega^{2}}{16}R_{\mathrm{DM}}$ $-\frac{R_{\mathrm{DM}}}{240}(5\omega^{3}-4k^{2}\omega)$
$\delta_{\gamma}$ $-\frac{2}{3}R_{\mathrm{DM}}\omega$ $\frac{\omega^{2}}{4}R_{\mathrm{DM}}$ $-\frac{R_{\mathrm{DM}}}{60}(5\omega^{3}-4k^{2}\omega)$
$\theta_{\gamma\mathrm{b}}/k$ $-\frac{1}{12}R_{\mathrm{DM}}k\omega$ $\frac{R_{\mathrm{DM}}(3R_{\mathrm{b}}-R_{\nu}+1)k\omega^{2}}{48(1-R_{\nu})}$
$\delta_{\nu_{\mathrm{eff}}}$ $-\frac{2}{3}R_{\mathrm{DM}}\omega$ $\frac{1}{4}R_{\mathrm{DM}}\omega^{2}$ $-\frac{R_{\mathrm{DM}}}{60}(5\omega^{3}-4k^{2}\omega)$
$\theta_{\nu_{\mathrm{eff}}}/k$ $-\frac{1}{12}R_{\mathrm{DM}}k\omega$ $\frac{R_{\mathrm{DM}}}{48}k\omega^{2}$
@@
$\delta_{\mathrm{b}}$ $-\frac{1}{2}R_{\mathrm{DM}}\omega$ $\frac{3\omega^{2}}{16}R_{\mathrm{DM}}$ $-\frac{R_{\mathrm{DM}}}{80}(5\omega^{3}-4k^{2}\omega)$
$\delta_{\mathrm{DM}}$ $1$ $-\frac{1}{2}R_{\mathrm{DM}}\omega$ $\frac{3\omega^{2}}{16}R_{\mathrm{DM}}$ $-\frac{R_{\mathrm{DM}}}{720}(45\omega^{3}-16k^{2}\omega)$
@@
To compare to the mixing angle formulation, we have to shift to $S_{\gamma\mathrm{b}}=0$ (cf. Section II). For this, the baryon and dark matter density perturbations need to be recasted here as well. It follows that $\delta_{\mathrm{b}}=\frac{3}{4}\delta_{\gamma}$ and, with the compensated matter perturbation condition $\delta\rho_{\mathrm{b}}+\delta\rho_{\mathrm{DM}}=0$, $\delta_{\mathrm{DM}}=\frac{3}{4}\frac{\rho_{\nu_{\mathrm{eff}}}}{\rho_{\gamma}}\frac{\rho_{\mathrm{b}}}{\rho_{\mathrm{DM}}}$. Retranslating into gauge-independent quantities yields
@@ Appendix B LSS and CMB observables
Figure 4: Comparison of the linear matter power spectrum today (redshift $z=0$) resulting from adiabatic, dark matter and neutrino isocurvature only as well as two admixed cases. For simplicity, we choose the scalar index $n=1$ and use the same amplitude $A_{s}$
@@
Large Scale Structure. It is well known that on scales that enter the horizon after matter-radiation equality, the curvature perturbation $\zeta_{\mathrm{mat}}$ corresponds to a constant Newtonian gravitational potential $\Phi$ and a matter perturbation $\delta_{\mathrm{m}}$ growing linearly in the scale factor as long as the contribution of dark energy to the total energy is negligible. When considering an adiabatic initial condition the curvature perturbation in matter domination is trivially related to the primordial one, $\zeta_{\mathrm{mat}}=\zeta_{\mathrm{p}}$ In the case of vanishing primordial curvature perturbation but with isocurvature of the form
@@
In the case of isocurvature fluctuations, the gravitational source terms, e.g. the Newtonian potential, vanish before horizon crossing. However, due to the difference in dynamics between neutrinos and the baryon-photon fluid, after horizon crossing, overdensities in the total radiation density build up and generate such terms. For the logarithmic growing solution to be present one therefore requires $S_{\gamma\nu}\neq 0$ Indeed we find that for the pure matter perturbation $\varphi=0$ the matter power spectrum falls proportional to $k^{-3}$ at large $k$, while for all other initial conditions eventually the logarithmically enhanced term proportional to $k^{-3}\log^{2}(k)$ comes to dominate as we have indicated in Fig. 4.
@@
The Newtonian potential $\Phi$ is given by $\Phi=3/5\ \zeta_{\mathrm{mat}}$ After the photons start free-streaming, they additionally get redshifted as they climb out of the gravitational potential. The observed temperature fluctuation is therefore given by
@@
Appendix C Comment on correlation with the adiabatic mode
In Fig. 6 we compare the posterior distribution as found by the MCMC projected onto the plane of isocurvature fraction $f_{\mathrm{iso}}$ and correlation with the adiabatic mode $c_{\zeta S}$ for compensated ($\varphi\approx 1.12$) and pure ($\varphi=\pi/2$) neutrino isocurvature. The slight improvement of the overall fit resulting in the reduction of $\chi^{2}$ when including pure neutrino isocurvature is visible here by the $1\sigma$ not overlapping with the $f_{\mathrm{iso}}=0$ axis as opposed to the case of compensated neutrino isocurvature.
Further we note that correlation with the adiabatic mode is disfavored in both cases. In the case of curvaton models predicting fully correlated ($c_{\zeta S}=\pm 1$) isocurvature [32, 33, 11, 12], the bounds on the isocurvature fraction are therefore significantly more stringent.

### v3

@@ III.2 Fit to CMB and LSS data
In Fig. 3 we show the dependence of the 95 percentile of $f_{\mathrm{iso}}^{2}=(P_{S_{\gamma\nu}}+P_{S_{\gamma\mathrm{DM}}})/P_{\zeta}|_{k_{\mathrm{piv}}}$ as a function of $\varphi$  in orange. For comparison the bounds found in Ref. [7] are indicated by green arrows. As the analysis of Ref. [7] allowed the isocurvature spectral index to vary, we compare our results to the most stringent one out of the three scales considered in Tab. 14 of Ref. [7], the one at $k_{\mathrm{low}}\,$. More specifically, we took the values $f_{\mathrm{iso}}^{2}=2.5\%$ and $7.4\%$ for the dark matter and compensated neutrino case, respectively, where for the latter one we accounted for the difference in normalization by multiplying with $(1.4)^{2}$. Our results agree with Ref. [7] to within 30%, where the discrepancies are most likely due to differences in analysis.
[…]
In Fig. 4 we compare the posterior distribution as found by the MCMC projected onto the plane of isocurvature fraction $f_{\mathrm{iso}}$ and correlation with the adiabatic mode $c_{\zeta S}$ for compensated ($\varphi\approx 1.12$) and pure ($\varphi=\pi/2$) neutrino isocurvature. We find that in the case of no correlation and pure neutrino isocurvature, a non-zero isocurvature fraction is preferred, while for the compensated neutrino isocurvature used by the Planck collaboration, this effect is not visible, emphasizing again the importance of mixed isocurvature modes.
Further we note that correlation with the adiabatic mode is disfavored in both cases, in agreement with Ref. [7]. In the case of curvaton models predicting fully correlated ($c_{\zeta S}=\pm 1$) isocurvature [45, 41, 35, 23], the bounds on the isocurvature fraction are therefore significantly more stringent.
@@ Acknowledgements.
We would like to thank J. Lesgourgues for insights and discussions on MontePython. The authors gratefully acknowledge the computing time granted on the supercomputer MOGON 2 at Johannes Gutenberg University Mainz (hpc.uni-mainz.de). PS and CG acknowledge support by the Cluster of Excellence “Precision Physics, Fundamental Interactions, and Structure of Matter” (PRISMA+ EXC 2118/1 and PRISMA++ EXC 2118/2) funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) within the German Excellence Strategy (Project No. 390831469).
@@ Appendix A Initial conditions in coordinates
$h$ $R_{\nu}\omega$ $\frac{3}{8}R_{\nu}\omega^{2}$ $\frac{R_{\nu}k^{2}\omega}{90}(3R_{\mathrm{b}}+4)-\frac{R_{\nu}\omega^{3}}{8}$
@@
$\frac{R_{\nu}(16R_{\nu}(3R_{\nu}+40)-175)k^{2}\omega}{360(2R_{\nu}+15)(4R_{\nu}+15)}$
$\delta_{\gamma}$ $\frac{4R_{\nu}}{3}$ $\frac{2R_{\nu}\omega\tau}{3}$ $\frac{1}{36}R_{\nu}\left(8k^{2}-9\omega^{2}\right)$
@@
$\frac{2R_{\nu}(R_{\mathrm{b}}(R_{\nu}-6))k^{2}\omega}{90(R_{\nu}-1)}$
$\theta_{\gamma\mathrm{b}}/k$ $\frac{R_{\nu}k}{3}$ $\frac{R_{\nu}R_{\mathrm{b}}k\omega}{4(R_{\nu}-1)+\frac{R_{\nu}k\omega}{12}}$
@@
$\frac{3R_{\nu}R_{\mathrm{b}}^{2}k\omega^{2}}{16(R_{\nu}-1)^{2}}$
@@
$\frac{R_{\nu}k\omega^{2}}{48}$
$\sigma_{\nu_{\mathrm{eff}}}$ $\frac{2(R_{\nu}-1)k^{2}}{12R_{\nu}+45}$ $\frac{(61-4R_{\nu})R_{\nu}k^{2}\omega}{18(2R_{\nu}+15)(4R_{\nu}+15)}$
$F_{3{\nu_{\mathrm{eff}}}}$ $\frac{4(R_{\nu}-1)k^{3}}{21(4R_{\nu}+15)}$
$\delta_{\mathrm{b}}$ $R_{\nu}$ $\frac{R_{\nu}\omega}{2}$ $\frac{1}{48}R_{\nu}\left(8k^{2}-9\omega^{2}\right)$
@@
$\frac{R_{\nu}R_{\mathrm{b}}(R_{\nu}-6)k^{2}\omega}{60(R_{\nu}-1)}$
$\delta_{\mathrm{DM}}$ $R_{\nu}$ $\frac{R_{\nu}\omega}{2}$ $\frac{3}{16}R_{\nu}\omega^{2}$ $\frac{R_{\nu}\omega^{3}}{16}-\frac{R_{\nu}k^{2}\omega(3R_{\mathrm{b}}+4)}{180}$
@@
$h$ $R_{\mathrm{DM}}\omega$ $\frac{3\omega^{2}}{8}R_{\mathrm{DM}}$ $\frac{R_{\mathrm{DM}}}{360}(45\omega^{3}-16k^{2}\omega)$
$\eta$ $\frac{1}{6}R_{\mathrm{DM}}\omega$ $\frac{\omega^{2}}{16}R_{\mathrm{DM}}$ $\frac{R_{\mathrm{DM}}}{240}(5\omega^{3}-4k^{2}\omega)$
$\delta_{\gamma}$ $\frac{2}{3}R_{\mathrm{DM}}\omega$ $\frac{\omega^{2}}{4}R_{\mathrm{DM}}$ $\frac{R_{\mathrm{DM}}}{60}(5\omega^{3}-4k^{2}\omega)$
$\theta_{\gamma\mathrm{b}}/k$ $\frac{1}{12}R_{\mathrm{DM}}k\omega$ $\frac{R_{\mathrm{DM}}(3R_{\mathrm{b}}-R_{\nu}+1)k\omega^{2}}{48(1-R_{\nu})}$
$\delta_{\nu_{\mathrm{eff}}}$ $\frac{2}{3}R_{\mathrm{DM}}\omega$ $\frac{1}{4}R_{\mathrm{DM}}\omega^{2}$ $\frac{R_{\mathrm{DM}}}{60}(5\omega^{3}-4k^{2}\omega)$
$\theta_{\nu_{\mathrm{eff}}}/k$ $\frac{1}{12}R_{\mathrm{DM}}k\omega$ $\frac{R_{\mathrm{DM}}}{48}k\omega^{2}$
@@
$\delta_{\mathrm{b}}$ $\frac{1}{2}R_{\mathrm{DM}}\omega$ $\frac{3\omega^{2}}{16}R_{\mathrm{DM}}$ $\frac{R_{\mathrm{DM}}}{80}(5\omega^{3}-4k^{2}\omega)$
$\delta_{\mathrm{DM}}$ 11 $\frac{1}{2}R_{\mathrm{DM}}\omega$ $\frac{3\omega^{2}}{16}R_{\mathrm{DM}}$ $\frac{R_{\mathrm{DM}}}{720}(45\omega^{3}-16k^{2}\omega)$
@@
To compare to the mixing angle formulation, we have to shift to $S_{\gamma\mathrm{b}}=0$ (cf. Section II). For this, the baryon and dark matter density perturbations need to be recast here as well. It follows that $\delta_{\mathrm{b}}=\frac{3}{4}\delta_{\gamma}$ and, with the compensated matter perturbation condition $\delta\rho_{\mathrm{b}}+\delta\rho_{\mathrm{DM}}=0$, $\delta_{\mathrm{DM}}=\frac{3}{4}\frac{\rho_{\nu_{\mathrm{eff}}}}{\rho_{\gamma}}\frac{\rho_{\mathrm{b}}}{\rho_{\mathrm{DM}}}$. Retranslating into gauge-independent quantities yields
@@ Appendix B LSS and CMB observables
Figure 5: Comparison of the linear matter power spectrum today (redshift $z=0$) resulting from adiabatic, dark matter and neutrino isocurvature only as well as two admixed cases. For simplicity, we choose the scalar index $n=1$ and use the same amplitude $A_{s}$ .
@@
Large Scale Structure. It is well known that on scales that enter the horizon after matter-radiation equality, the curvature perturbation $\zeta_{\mathrm{mat}}$ corresponds to a constant Newtonian gravitational potential $\Phi$ and a matter perturbation $\delta_{\mathrm{m}}$ growing linearly in the scale factor as long as the contribution of dark energy to the total energy is negligible. When considering an adiabatic initial condition the curvature perturbation in matter domination is trivially related to the primordial one, $\zeta_{\mathrm{mat}}=\zeta_{\mathrm{p}}$ . In the case of vanishing primordial curvature perturbation but with isocurvature of the form
@@
In the case of isocurvature fluctuations, the gravitational source terms, e.g. the Newtonian potential, vanish before horizon crossing. However, due to the difference in dynamics between neutrinos and the baryon-photon fluid, after horizon crossing, overdensities in the total radiation density build up and generate such terms. For the logarithmic growing solution to be present one therefore requires $S_{\gamma\nu}\neq 0$ . Indeed we find that for the pure matter perturbation $\varphi=0$ the matter power spectrum falls proportional to $k^{-3}$ at large $k$, while for all other initial conditions eventually the logarithmically enhanced term proportional to $k^{-3}\log^{2}(k)$ comes to dominate as we have indicated in Fig. 5.
@@
The Newtonian potential $\Phi$ is given by $\Phi=3/5\ \zeta_{\mathrm{mat}}$ . After the photons start free-streaming, they additionally get redshifted as they climb out of the gravitational potential. The observed temperature fluctuation is therefore given by